What could be the system of linear equations for this problem if i want to use elimination or substitution?
while on vacation, kevin went for a swim in a nearby lake. swimming against the current, it took him 8 minutes to swim 200 meters. swimming back to shore with the current took half as long. what is kevins average swimming speed and speed of lake's current? i'd like to use variables x and y.

Answer by ankor@dixie-net.com(22740)

What could be the system of linear equations for this problem if i want to use elimination or substitution?
:
While on vacation, kevin went for a swim in a nearby lake. Swimming against the current, it took him 8 minutes to swim 200 meters. swimming back to shore with the current took half as long. what is Kevins average swimming speed, and speed of the lake's current? I'd like to use variables x and y.
:
Let x = swimmer's speed; y = lake current speed
:
speed against = (x-y) in meters/min
speed with = (x+y)
:
Write two distance equations: Time * Speed = Dist
8(x-y) = 200; simplify, divide eq by 8, resulting in:
4(x+y) = 200; simplify, divide eq by 4, resulting in:
:
x - y = 25
x + y = 50
------------ add
2x + 0 = 75
x = 75/2
x = 37.5 m/min swim speed
:
Find y:
37.5 - y = 25
y = 12.5 m/min is the lake current
:
Check out solutions:
8(37.5 - 12.5) = 200
;
4(37.5 + 12.5) = 200 also
: